Find f(-2),f(0) and f(4) for the following function.
f(x)=-3x
f(-2)=
f(0)=
f(4)=

Answer by Shin123(626)

f(-2)=-3*-2=6
f(0)=-3*0=0
f(4)=-3*4=12.